OpenStudy (nortexplay):

Which information is not necessary to solve this problem? Bruce works for 8 hours a day. He is paid after every 5 days of work. He has been working at his job for 60 days. He makes $600 each pay check. How much does Bruce get paid each hour? A. Bruce works 8 hours a day. B. Bruce gets paid every 5 days. C. Bruce has been at his job for 60 days. D. Bruce makes $600 a paycheck.

OpenStudy (nortexplay):

I think it's b

OpenStudy (nortexplay):

Cause it says which answer is not necessary. it's b right?

OpenStudy (first_world_problems):

could be c

OpenStudy (nortexplay):

why?

OpenStudy (first_world_problems):

but idk

OpenStudy (nortexplay):

.-.

OpenStudy (nortexplay):

it says Which information is not necessary to solve this problem

OpenStudy (nortexplay):

b seems none needed

OpenStudy (nortexplay):

idkdkdkdkk

OpenStudy (first_world_problems):

@agent0smith help

OpenStudy (first_world_problems):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`Bruce works for 8 hours a day. He is paid after every 5 days of work. He has been working at his job for 60 days. He makes $600 each pay check. ` He works 8 hrs a day for 5 days, so he works 8*5 = 40 hours a week (excluding weekends) If he gets paid $600 every 5 days, or every week, then he earns (600 dollars)/(40 hours) = (600/40)(dollars/hour) = 15 dollars per hour Notice how I didn't use the fact that he has been working at his job for 60 days. It doesn't matter. He still gets paid the same hourly wage whether he has been working for 6 days or 60 days.

OpenStudy (nortexplay):

so c :(

jimthompson5910 (jim_thompson5910):

yes it's C

OpenStudy (first_world_problems):

Je pété ...